# SaVe-Tag: Semantic-aware Vicinal Risk Minimization for Long-Tailed Text-Attributed Graphs

### v3

@@ Appendix B Implementation Details
B.1 Code Implementation
Algorithm
Our codes are provided in the following link:
@@ B.1 Code Implementation
https://github.com/LWang-Laura/SaVe-TAG
The pseudocode of SaVe-Tag  is provided in Algorithm 1.
1
2
Input: Attributed graph $\mathcal{G}=(\mathcal{V},\mathcal{T},\mathcal{E})$ with texts $\mathcal{X}$ and labels $\mathcal{Y}$;
vicinal order $k$; temperature–controlled LLM $\mathcal{L}_{\tau}$;
frozen encoder $\phi$; confidence network $\kappa_{\omega}$;
edge budget $k_{\text{edge}}$; confidence threshold $\tau$
Output: Augmented graph $\mathcal{G}^{\prime}=(\mathcal{V}^{\prime},\mathcal{T}^{\prime},\mathcal{E}^{\prime})$; trained classifier $f_{\theta}$
3 Initialize $\bar{\mathcal{T}}$, $\bar{\mathcal{V}}$, $\bar{\mathcal{E}}$ as empty sets
4 Identify tail classes: $\mathcal{C}_{t}\leftarrow\{c\mid\text{freq}(c)<\text{median freq}\}$
5 $\mathcal{V}_{t}\leftarrow\{v\in\mathcal{V}\mid y_{v}\in\mathcal{C}_{t}\}$
6
7
8 Function DataAugment($v,u$):
9       $\hat{t}\leftarrow\mathcal{L}_{\tau}\bigl{(}t_{v},t_{u}\,\big{|}\,y_{v}\bigr{)}$
10       Initialize new node $\hat{v}$ with text $\hat{t}$ and label $y_{\hat{v}}=y_{v}$
11       return $\hat{v}$
12
13
// Generate vicinal twins and interpolate with the LLM
14 for $v\in\mathcal{V}_{t}$ do
15       $\mathcal{N}^{(\mathrm{tail})}_{k}(v)\leftarrow\{u\in\mathcal{V}_{t}\mid u\in\mathcal{N}_{k}(v)\}$
16       for $u\in\mathcal{N}^{(\mathrm{tail})}_{k}(v)$ do
17             $\hat{v}\leftarrow\texttt{DataAugment}(v,u)$
18             Append $\hat{v}$ to $\bar{\mathcal{V}}$, its text $\hat{t}$ to $\bar{\mathcal{T}}$
19
20
21
22
// Pre-train confidence function on original graph
23 $\kappa_{\omega}\leftarrow\text{train}(\{z_{v}=\phi(t_{v})\}_{v\in\mathcal{V}})$
24
25
// Edge assignment guided by confidence
26 for $\hat{v}\in\bar{\mathcal{V}}$ do
27       $z_{\hat{v}}\leftarrow\phi(\hat{t}_{\hat{v}})$
28       $s_{u}\leftarrow\kappa_{\omega}(z_{u})\langle z_{\hat{v}},z_{u}\rangle$ for $u\in\mathcal{V}$
29       $\mathcal{N}_{k_{\text{edge}}}(\hat{v})\leftarrow\arg\mathrm{top}_{k_{\text{edge}}}s_{u}$
30       if $\max_{u}s_{u}\geq\tau$ then
31             ¯¯$\bar{\mathcal{E}}\leftarrow\bar{\mathcal{E}}\cup\{(\hat{v},u)\mid u\in\mathcal{N}_{k_{\text{edge}}}(\hat{v})\}$
32
33      else
34             pass;
35
36
37
38Form augmented graph:
39 $\mathcal{V}^{\prime}\leftarrow\mathcal{V}\cup\bar{\mathcal{V}}$,
40 $\mathcal{T}^{\prime}\leftarrow\mathcal{T}\cup\bar{\mathcal{T}}$,
41 $\mathcal{E}^{\prime}\leftarrow\mathcal{E}\cup\bar{\mathcal{E}}$
42
43Train GNN $f_{\theta}$ on $\mathcal{G}^{\prime}$
44
Result: $\mathcal{G}^{\prime},\,f_{\theta}$
45
Algorithm 1 SaVe-Tag
Computing Environment and Resources
We implement our model under PyG [37] and Sentence Transformer [38] modules. Experiments are conducted on a NVIDIA GeForce RTX 3090 and the OS was Ubuntu 22.04.4 LTS with 128GB RAM.
B.2 Hyperparameters
The detailed hyperparameters of our setups for reproducibility are listed below:
  • Text Encoding: We utilize the Sentence Transformer (all-MiniLM-L6-v2) [38], for its efficiency in semantic comprehension. This model strikes an excellent balance between being lightweight and delivering strong performance. A comparison with other embeddings is provided in Figure 10.
  • Node Classification: By default, we use GCN as our model, consisting of 2 hidden layers with 64 neurons each. The dropout rate is 0.5 and the model is trained for 1000 epochs with a learning rate of 0.01 (same configuration as we switch to SAGE when conducting experiments in Table 4).
  • Confidence Function: We employ MLP models composed of 1 hidden layer with 256 neurons and set with a dropout rate of 0, which are trained for 1000 epochs with a learning rate of 0.001.
  • Text Generation: By default, we leverage pre-trained Llama3-8B-Instruct from Meta for text generation, configured with bfloat16 and default parameters.
  • Data Augmentation: Due to the limited number of training nodes in our low-labeled, imbalanced settings, we selected $k=3$ for the k-nearest neighbors in our LLM-based methods.
  • Edge generation: For small datesets (Cora, PubMed, and Citeseer), we set $k=|\mathcal{V}|\times 20$ and add the top $k$ highest prediction score edges to the graph. For larger datasets (Photo, Computer, and Children), we increase $k$ to $|\mathcal{V}|\times 40$ correspondingly.
B.3 Datasets
In this part, we include brief descriptions of each dataset. For Cora and PubMed, we downloaded the preprocessed datasets from https://github.com/CurryTang/Graph-LLM [39]. As to Citeseer, Photo, Children, and Computer datasets, the preprocessed graph data are downloaded from https://github.com/CurryTang/TSGFM [40] and their detailed descriptions can be found in https://github.com/sktsherlock/TAG-Benchmark [30]. The category names for each dataset are listed below, with their indices corresponding to the numeric labels in the dataset.
  • Cora: [‘Rule Learning’, ‘Neural Networks’, ‘Case Based’, ‘Genetic Algorithms’, ‘Theory’, ‘Reinforcement Learning’, ‘Probabilistic Methods’]
  • PubMed: [‘Diabetes Mellitus, Experimental’, ‘Diabetes Mellitus Type 1’, ‘Diabetes Mellitus Type 2’]
  • Citeseer:[‘Agents’, ‘ML(Machine Learning)’, ‘IR(Information retrieval)’, ’DB(Database)’, ‘HCI (Human-computer Interaction)’, ‘AI(Artificial Intelligence)’]
  • Photo:[ ‘Video Surveillance’,"Accessories’,‘Binoculars & Scopes’, ‘Video’, ‘Lighting & Studio’, ‘Bags & Cases’,‘Tripods & Monopods’, ‘Flashes’, ‘Digital Cameras’, ‘Film Photography’, ‘Lenses’, ‘Underwater Photography’]
  • Children: [‘Literature & Fiction’, ‘Animals’, ‘Growing Up & Facts of Life’, ‘Humor’, ‘Cars Trains & Things That Go’, ‘Fairy Tales Folk Tales & Myths’, ‘Activities Crafts & Games’, ‘Science Fiction & Fantasy’, ‘Classics’, ‘Mysteries & Detectives’, ‘Action & Adventure’, ‘Geography & Cultures’, ‘Education & Reference’, ‘Arts Music & Photography’, ‘Holidays & Celebrations’, ‘Science Nature & How It Works’, ‘Early Learning’, ‘Biographies’, ‘History’, “Children’s Cookbooks", ‘Religions’, ‘Sports & Outdoors’, ’Comics & Graphic Novels’, ’Computers & Technology’]
  • Computer: [‘Computer Accessories & Peripherals’, ‘Tablet Accessories’, ‘Laptop Accessories’, ‘Computers & Tablets’, ‘Computer Components’, ‘Data Storage’, ‘Networking Products’, ‘Monitors’, ‘Servers’, ‘Tablet Replacement Part’]
B.4 Baselines Details
@@
  • Zero-shot [10, 11]: Utilizing a pre-trained LLM to generate textual attributes for a given target label without providing specific examples.
  • Few-shot [10, 11]: Prompting an LLM to generate textual attributes for a given target label, providing examples for in-context learning.
@@
  • Oversampling [32]: Repeatedly duplicates node embeddings that belong to the minority classes.
  • SMOTE [7]: Generating synthetic nodes by interpolating tail nodes with neighbors and assigning edges by copying their neighbors’ edges.
  • Embed-SMOTE [33]: Applying SMOTE on hidden layer embeddings rather than the input features.
  • MixupForGraph [9]: Synthesizing additional data for minority classes by interpolating randomly paired hidden representation as well as their labels
  • GraphSMOTE [8]: Interpolating minority nodes with their nearest neighbors and generating new edges using a co-trained link predictor, with two variants denoted by subscript T and O, depending on whether the predicted edges are discrete or continuous.
  • LTE4G [3]: Clustering nodes by class and degree, training subset-specific experts, and distilling them into head and tail students for classification.
  • HierTail [2]: Extracting shared class information via hierarchical task grouping and balancing head–tail gradient contributions.
@@
The description of the three variants (SaVe-TagS, SaVe-TagM, SaVe-TagO) are listed in Section 5.1.3. A case study illustration is provided in Figure 13.
Figure 13: A case study illustrating three variants our LLM-based interpolation: (a) SaVe-TagO, (b) SaVe-TagS, and (c) SaVe-TagM.
@@
User: Give me the first {Text} from {Dataset} with topic [$\mathcal{C}_{1}$].
@@
User: Give me the second {Text} from {Dataset} with topic [$\mathcal{C}_{2}$].
@@
User: Give me the third {Text} from {Dataset} with topic [$\mathcal{C}_{2}$]. It should be more similar to the first {Text} and less similar to the second {Text}.
@@
Table 5: Prompt template for SaVe-TagS and SaVe-TagM (${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathcal{C}_{1}}={\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathcal{C}_{2}}$ for SaVe-TagO).
System: You are a helpful AI assistant for generating {Task} from {Dataset}, where each {Task} follows the format <START>{Format}<END>.
User: Give me the first {Task} from {Dataset} with topic [$\mathcal{C}_{1}$].
@@
User: Give me the second {Task} from {Dataset} with topic [$\mathcal{C}_{1}$]. It should be more similar to the first {Task}.
@@
Table 6: Prompt template for SaVe-TagO across all datasets.
@@
Table 7: Input parameters for different datasets in the prompt templates.
B.5 Prompt Design
The prompt template for SaVe-TagO is shown in Table 6, while SaVe-TagS and SaVe-TagM share a common template in Table 5, with the constraint of shared class ( ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathcal{C}_{1}}={\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathcal{C}_{2}}$ ) when using SaVe-TagS. Both templates are dataset-specific, with their parameters detailed in Table 7.
@@ Appendix C Efficiency and Scalability Analysis
Name # Nodes # Edges # Class # Tail Ave Text Len # Batch Prompts/Batch Time/Prompt Cora 2,708 10,858 7 5 891 20 4 3.06s PubMed 19,717 88,670 3 2 1649 8 4 4.51s Citeseer 3186 8450 6 4 1022 16 4 3.50s Photo 48,362 500,939 12 8 804 64 2 4.60s Computer 87,229 721,081 10 6 499 24 2 4.09s Children 76,875 1,554,578 24 15 1255 240 1 7.21s
Table 8: Dataset Details. We recorded the following statistics on all datasets: the number of nodes, edges, and classes; the number of tail (minority) classes; the average text length of node attributes; the total number of batches; the number of prompts per batch (batch size); and the average time per prompt (generated entry) during generation.
C.1 Time Costs
For scalability analysis, we report the statistics regarding the size of the graph (number of edges, nodes, average length of text attributes, etc.) in the Table 8.
@@ C.1 Time Costs
C.2 Adaptability for Small Language Models
Figure 14: The F1 score and generation time per entry of LLM variants across different datasets implementing SaVe-TagS.
@@ C.2 Adaptability for Small Language Models
Figure 14 shows the performance and efficiency of SaVe-TagS using language models with different sizes, including Llama-3-8B-Instruct, Llama-3.2-1B-Instruct, and Llama-3.2-3B-Instruct. As expected, smaller models require less inference time for text generation. Notably, the F1 score remains relatively stable across model sizes, highlighting the adaptability of our method to lightweight, open-source LLMs—a promising direction for reducing generation costs.
Appendix D Related Works
Vicinal Risk Minimization (VRM)
VRM [35] is a principle in statistical learning theory that extends the traditional Empirical Risk Minimization (ERM) framework [5]. Introduced by Chapelle et al., VRM proposes estimating the expected risk by considering a vicinity distribution around each training sample, rather than relying solely on the empirical distribution. This approach aims to improve generalization by incorporating local neighborhoods of data points into the learning process. VRM has been foundational in developing techniques like Mixup [6] and SMOTE [7], which generate synthetic training examples through linear interpolations of existing data points, thereby enhancing model generalization capabilities.
  • Mixup [6]: Given $(x_{i},y_{i}),(x_{j},y_{j})$, draw $\lambda\sim\mathrm{Beta}(\alpha,\alpha)$ and set
    $\tilde{x}=\lambda x_{i}+(1-\lambda)x_{j},\quad\tilde{y}=\lambda y_{i}+(1-\lambda)y_{j}.$
  • SMOTE [7]: For minority sample $x_{i}$, choose a neighbor $x_{k}$, draw $\lambda\sim U(0,1)$ and define
    $\tilde{x}=x_{i}+\lambda\,(x_{k}-x_{i}),\quad\tilde{y}=y_{i}=y_{j}.$
Long-tail Graph Learning
In many real-world graph-structured datasets, the distribution of node classes follows a long-tailed pattern, where a few classes (head classes) have abundant samples, while many others (tail classes) have scarce representations. This imbalance poses significant challenges for graph neural networks (GNNs), which may become biased towards
  • Hierarchical Task Grouping: Formulating long-tailed classification as a multi-task learning problem, where each class prediction is treated as a separate task. This approach allows for shared representations among related tasks, improving performance on tail classes [2].
  • Expert Models: Training specialized models (experts) for different subsets of classes or nodes, such as the LTE4G [3], which considers both class and degree long-tailedness in graphs.
LLM Data Augmentation
Large Language Models (LLMs) have revolutionized data augmentation strategies, especially in natural language processing tasks. By generating synthetic data that mirrors the distribution and characteristics of real data, LLMs can enhance model training in low-resource settings. Key approaches include:
  • Instruction-based Augmentation: Empowering LLMs with task-specific instructions to generate diverse and contextually relevant data [36].
  • Zero-shot or Few-shot Augmentation: Enabling LLMs to generate synthetic examples with minimal or no task-specific labels via in-context learning[10].
Appendix E Ethics and Broader Impact
SaVe-Tag  LLMs to generate synthetic text for augmenting long-tailed node classification on text-attributed graphs. By mitigating class imbalance, it improves classification performance on widely used benchmark datasets such as Cora, Citeseer, and PubMed. The use of lightweight, open-source LLMs and one-time data augmentation helps reduce computational cost, potentially enabling broader access to graph learning techniques in low-resource settings.
Our study is conducted entirely on publicly available, anonymous datasets, and does not involve any personally identifiable or sensitive information. The generated synthetic data is stored in files and used exclusively for research purposes. While our method does not raise significant ethical concerns in its current scope, we encourage practitioners to clearly label synthetic data in downstream applications to avoid confusion between real and generated content.
Appendix F Detailed Proof
F.1 Proof of Theorem 4.4 (Margin lower bound)
Notation recap.
$\gamma(x)$ is the logit margin of a point $x$, $\gamma_{\min}(S)=\min_{x\in S}\gamma(x)$, and $\gamma_{0}:=\gamma_{\min}(\mathcal{D})$ is the smallest margin on the original training set $\mathcal{D}$. After augmentation, we have $\tilde{\mathcal{D}}\;=\;\mathcal{D}\cup\widehat{\mathcal{D}}_{\mathrm{bd}}\cup\widehat{\mathcal{D}}_{\mathrm{in}},$ where $\widehat{\mathcal{D}}_{\mathrm{bd}}$ (resp. $\widehat{\mathcal{D}}_{\mathrm{in}}$) holds all synthetic boundary (resp. interior) samples and $\mathrm{BCR}=\frac{|\widehat{\mathcal{D}}_{\mathrm{bd}}|}{|\widehat{\mathcal{D}}_{\mathrm{bd}}|+|\widehat{\mathcal{D}}_{\mathrm{in}}|}\in[0,1].$
@@ Appendix F Detailed Proof
    Conservative ordering. We choose $\delta\!\geq\!\gamma_{0}$. This turns A2. into a trivial worst-case cap (it can never be violated) and lets the algebra below produce a bona-fide lower bound—that is, the right-hand side never exceeds the true minimum margin.
@@
Case 1: $\boldsymbol{m_{\mathrm{bd}}=0}$ (BCR $=0$). The augmented set contains only $\mathcal{D}\cup\widehat{\mathcal{D}}_{\mathrm{in}}$. By Assumption A1. every margin is still $\geq\gamma_{0}$, hence $\gamma_{\min}(\tilde{\mathcal{D}})\geq\gamma_{0}.$ The bound in the statement reduces to $\gamma_{0}-\delta(1-0)=\gamma_{0},$ so the inequality is tight in this corner case.
Case 2: $\boldsymbol{m_{\mathrm{bd}}>0}$ (BCR $>0$). Now at least one boundary sample is present. By A2. its margin is $\geq-\delta$ and, because $\delta\geq\gamma_{0}$ (A3.), also $\geq\gamma_{0}-\delta$. Combining with Assumption A1. for all other points gives
@@
Because BCR$\in(0,1]$, $\gamma_{0}-\delta(1-\mathrm{BCR})\;\leq\;\gamma_{0}-\delta,$ and the claimed inequality again holds.
@@
F.2 Proof of Theorem 4.5 (On-Manifold Vicinal Risk Theorem)
Notation recap.
For each original training pair $(t_{i},c_{i})$:
  • $\mathcal{V}_{\mathcal{L}_{\tau}}(t_{i})$ is the vicinal distribution produced by LLM-based interpolation.
  • $\hat{t}\sim\mathcal{V}_{\mathcal{L}_{\tau}}(t_{i})$ is one synthetic neighbor.
  • $\phi:\mathcal{T}\!\to\!\mathbb{R}^{d^{\prime}}$ is the frozen encoder; $\mathcal{M}_{c_{i}}=\{\phi(t):y=c_{i}\}$ is the class-$c_{i}$ manifold.
  • $\mathbf{L}(\cdot,\cdot)$ is the loss (cross-entropy, hinge, …).
  • $f_{\theta}$ is the classifier that we are analyzing.
@@
    Finite loss.$\mathbf{L}(f_{\theta}(\phi(\hat{t})),c_{i})$ is integrable so that all expectations are well defined. [^2] A bounded loss (e.g. cross-entropy with label smoothing) or a sub-Gaussian surrogate satisfies this automatically.
@@
Fix $i$ temporarily. Apply A1. and the law of total expectation [41]:
@@
$R_{\text{on}}:=\frac{1}{n}\sum_{i=1}^{n}\mathbb{E}\!\bigl{[}\mathbf{L}(\cdot)\mid\phi(\hat{t})\in\mathcal{M}_{c_{i}}\bigr{]},\qquad R_{\text{off}}:=\frac{1}{n}\sum_{i=1}^{n}\mathbb{E}\!\bigl{[}\mathbf{L}(\cdot)\mid\phi(\hat{t})\notin\mathcal{M}_{c_{i}}\bigr{]}.$
Summing the displayed equality over $i=1,\dots,n$ and using A2. yields
@@
F.3 Proof of Theorem 4.6 (Boundary-Coverage $\Rightarrow$ Vicinal-Risk Reduction)
Notation recap.
  • $\tilde{\mathcal{D}}=\mathcal{D}\cup\widehat{\mathcal{D}}$ is the augmented training set; $\widehat{\mathcal{D}}$ contains $m$ synthetic points.
  • $\widehat{\mathcal{D}}_{\mathrm{bd}}$ / $\widehat{\mathcal{D}}_{\mathrm{in}}$ synthetic boundary / interior subsets; $\mathrm{BCR}:=|\widehat{\mathcal{D}}_{\mathrm{bd}}|/m$.
  • $\gamma(\hat{x})$ logit margin of $\hat{x}$; $\gamma_{0}:=\min_{x\in\mathcal{D}}\gamma(x)$.
  • $\delta$ upper bound on $|\gamma(\hat{x})|$ for $\hat{x}\in\widehat{\mathcal{D}}_{\mathrm{bd}}$ (Theorem 4.4).
  • $\mathbf{L}$ supervised loss and $L$ is the Lipschitz constant of $\mathbf{L}$ in its first (logit) argument.
  • $f_{\theta}^{\mathrm{orig}}$ model trained on $\mathcal{D}$; $f_{\theta}^{\mathrm{aug}}$ model after VRM training on $\tilde{\mathcal{D}}$.
@@
    Lipschitz loss. $\mathbf{L}(z_{1},y)-\mathbf{L}(z_{2},y)\leq L\|z_{1}-z_{2}\|_{2}$  [42].
@@
    Non-decreasing margins. VRM retraining never reduces the margin of a synthetic point: $\gamma_{\,\mathrm{aug}}(\hat{x})\geq\gamma_{\,\mathrm{orig}}(\hat{x})$.
@@
(Last three assumptions are justified in Appendix F.1.)
@@
For any synthetic $(\hat{x},\hat{y})$, combine A1. + A2. to obtain
$\mathbf{L}\bigl{(}f_{\theta}^{\mathrm{aug}}(\hat{x}),\hat{y}\bigr{)}-\mathbf{L}\bigl{(}f_{\theta}^{\mathrm{orig}}(\hat{x}),\hat{y}\bigr{)}\;\leq\;-L\,\gamma(\hat{x}).$
@@
Sum (F.3) over the two synthetic subsets and divide by $m$:
$R_{vrm}\!\bigl{(}f_{\theta}^{\mathrm{aug}}\bigr{)}-R_{vrm}\!\bigl{(}f_{\theta}^{\mathrm{orig}}\bigr{)}$ $\leq-\frac{L}{m}\sum_{\hat{x}\in\widehat{\mathcal{D}}_{\mathrm{bd}}}\!\!\gamma(\hat{x})-\frac{L}{m}\sum_{\hat{x}\in\widehat{\mathcal{D}}_{\mathrm{in}}}\!\!\gamma(\hat{x})$
A3.,A4.${\leq}-L\,\delta\,\frac{|\widehat{\mathcal{D}}_{\mathrm{bd}}|}{m}\;-\;L\,\gamma_{0}\,\frac{|\widehat{\mathcal{D}}_{\mathrm{in}}|}{m}.$
@@
Choose $\eta=\delta/\gamma_{0}\;(0<\eta<1)$ and rewrite (F.3) as
@@
F.4 Proof of Theorem 4.8 (Pulling Well-Aligned Nodes)
Notation.
  • $h^{(\ell)}_{v}$ – representation of node $v$ after $\ell$ GNN layers.
  • $\mathcal{M}_{c}$ – representation manifold of class $c$ in $\mathbb{R}^{p}$.
  • $d_{\ell}:=\operatorname{dist}\bigl{(}h^{(\ell)}_{\hat{v}},\mathcal{M}_{c}\bigr{)}$ – Euclidean distance of synthetic node $\hat{v}$ to $\mathcal{M}_{c}$.
  • $\mathbf{W}$ – linear map inside the GNN layer; $\alpha\in(0,1)$, $\beta_{\hat{v}u}\geq 0$ with $\sum_{u\in\mathcal{N}^{\prime}(\hat{v})}\beta_{\hat{v}u}=1$.
@@
    Well-aligned neighbors:$\operatorname{dist}\!\bigl{(}h^{(\ell)}_{u},\mathcal{M}_{c}\bigr{)}\leq\varepsilon$ for every $u\in\mathcal{N}^{\prime}(\hat{v})$.
@@
    Convexity of weights:$\alpha\in(0,1)$ and $\{\beta_{\hat{v}u}\}$ form a convex combination.

### v5

@@ Appendix B Implementation Details
Code Implementation. Our model uses PyG (Fey and Lenssen, 2019) and Sentence Transformer (Reimers and Gurevych, 2019). Experiments were conducted on Ubuntu 22.04.4 LTS with 128GB RAM and NVIDIA GeForce RTX 3090.
@@
Hyperparameters. Details of our setups for reproducibility can be found at https://github.com/LWang-Laura/SaVe-TAG.
Datasets. We obtained Cora and PubMed from (Chen et al., 2024a); Citeseer, Photo, Children, and Computer datasets from (Chen et al., 2024b) with descriptions in (Yan et al., 2023).
Baselines Details.
@@
  • Zero-shot (Li et al., 2023; Yu et al., 2024): Utilizing a pre-trained LLM to generate textual attributes for a given target label without providing examples.
  • Few-shot (Li et al., 2023; Yu et al., 2024): Prompting an LLM to generate text for a given target label, providing examples for in-context learning.
@@
  • Oversampling (Chawla, 2003): Repeatedly duplicates minority/tail nodes.
  • SMOTE (Chawla et al., 2002): Synthetic nodes interpolate tail nodes with neighbors and assigning edges by copying their neighbors’ edges.
  • Embed-SMOTE (Ando and Huang, 2017): Applying SMOTE on hidden layer embeddings rather than the input features.
  • MixupForGraph (Wang et al., 2021a): Synthesizing data for minority classes by interpolating random pairs of hidden representation and labels.
  • GraphSMOTE (Zhao et al., 2021): Interpolating minority nodes with nearest neighbors and generating new edges via a co-trained link predictor; two variants T and O, for discrete or continuous edges
  • LTE4G (Yun et al., 2022): Clustering nodes by class and degree, training subset experts, and then distilling them for classification.
  • HierTail (Wang et al., 2024): Extracting shared class information via hierarchical task grouping and balancing head–tail gradient contributions.
@@
The description of the three variants are listed in Section 4.1.3 and a case study is provided in Figure 9.
Figure 9. A case study illustrating our three variants:
(a) SaVe-TAGO, (b) SaVe-TAGS, and (c) SaVe-TAGM.
B.1. Prompt Design
The prompt template for SaVe-TAGS and SaVe-TAGM share a common template in Table 6; SaVe-TAGO is shown in Table 7. Templates are dataset-specific, with parameters detailed in Table 8.
Table 6. Prompt template for SaVe-TAGS and SaVe-TAGM (${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathcal{C}_{1}}={\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathcal{C}_{2}}$ for SaVe-TAGO).
@@ B.1. Prompt Design
User: Give me the first {Text} from {Dataset} with topic $\mathcal{C}_{1}$].
@@
User: Give me the second {Text} from {Dataset} with topic $\mathcal{C}_{2}$].
@@
User: Give me the third {Text} from {Dataset} with topic $\mathcal{C}_{2}$]. It should be more similar to the first {Text} and less similar to the second {Text}.
@@
Table 7. Prompt template for SaVe-TAGO across all datasets.
User: Give me the first {Text} from {Dataset} with topic $\mathcal{C}_{1}$].
@@
User: Give me the second {Text} from {Dataset} with topic $\mathcal{C}_{1}$]. It should be more similar to the first {Text}.
@@
Table 8. Prompt template input parameters for each dataset.
@@
Table 9. Dataset Details.
Name #Nodes #Edges #Class #Tail Len #Batch BZ Time Cora 2,708 10,858 7 5 891 20 4 3.06s PubMed 19,717 88,670 3 2 1649 8 4 4.51s Citeseer 3186 8450 6 4 1022 16 4 3.50s Photo 48,362 500,939 12 8 804 64 2 4.60s Computer 87,229 721,081 10 6 499 24 2 4.09s Children 76,875 1,554,578 24 15 1255 240 1 7.21s
  • Len: the avg. text length of node attributes; # Batch: the total number of batches; BZ (Batch Size): the number of prompts per batch; Time: the avg. time per prompt (generated entry) during generation.
Table 10. Runtime and memory usage across methods.
Cora Citeseer PubMed
Method Time RAM GPU Time RAM GPU Time RAM GPU
LTE4G 120.4 1.9 630 119.4 1.6 838 OOM OOM OOM
HierTail 67.2 1.5 630 49.8 1.6 838 OOM OOM OOM
SaVe-TAG 10.3 1.7 738 11.4 1.7 738 24.9 1.8 954
  • Note that Time is in seconds, RAM in GB, and GPU in MB.
@@ Appendix C Efficiency and Scalability Analysis
C.1. Time Costs
First, we report the statistics regarding the graph sizes in Table 9.
For scalability analysis and examine efficiency, we measure post-augmentation runtime of our method, and report peak RAM, peak GPU usage, and wall-clock time for each run. We compare SaVe-TAG with SOTA baselines LTE4G and HierTail. Notably, both LTE4G and HierTail encountered OOM issues on PubMed (We manually optimized their code to obtain the results reported in Table 10).
@@ C.1. Time Costs
C.2. Adaptability for Different LLM Variants
@@
Figure 10 shows the performance and efficiency of SaVe-TAGS using LLMs with different sizes, including Llama-3-8B-Instruct, Llama-3.2-1B-Instruct, and Llama-3.2-3B-Instruct. As expected, smaller models require less inference time for text generation. Notably, the F1 score remains relatively stable across model sizes, highlighting the adaptability of our method to lightweight, open-source LLMs—a promising direction for reducing generation costs.
Figure 10. The F1 score and generation time per entry of LLM variants across different datasets implementing SaVe-TAGS.
Appendix D Detailed Proofs
D.1. Proof of Theorem 3.2
Manifold-Preserving Class-Consistent Generation.
Assumptions for this theorem.
  1. A1.
[…]
  2. A2.
[…]
  3. A3.
    Measurability.AA is measurable so that $p_{c}(A)$ and $q_{c}(A)$ are well defined. [^1] This holds automatically when $\phi$ is measurable and $\mathcal{M}_{c}$ is Borel.
Proof.
By Assumption A2. and the definition of total variation, $q_{c}(A)\;\geq\;p_{c}(A)-D_{\mathrm{TV}}(p_{c},q_{c}).$ Applying Assumption A1. yields $q_{c}(A)\;\geq\;(1-\delta)-\epsilon\;=\;1-(\delta+\epsilon).$ By the definition of $A$ and $q_{c}$, $q_{c}(A)=\Pr_{\hat{t}\sim\mathcal{L}_{\tau}(\cdot\mid c)}\!\big[\phi(\hat{t})\in\mathcal{M}_{c}\big]$,
which proves $\Pr\big[\phi(\hat{t})\in\mathcal{M}_{c}\big]\;\geq\;1-(\delta+\epsilon).\quad$
D.2. Proof of Theorem 3.5
Margin Lower Bound.
@@ Appendix D Detailed Proofs
    Conservative ordering. We choose $\delta\!\geq\!\gamma_{0}$. This turns A2. into a trivial worst-case cap (it can never be violated) and lets the algebra below produce a bona-fide lower bound—that is, the right‐hand side never exceeds the true minimum margin.
@@
Case 1: $\boldsymbol{m_{\mathrm{bd}}=0}$ (BCR =0=0). The augmented set contains only $\mathcal{D}\cup\widehat{\mathcal{D}}_{\mathrm{in}}$. By Assumption A1. every margin is still $\geq\gamma_{0}$, hence $\gamma_{\min}(\tilde{\mathcal{D}})\geq\gamma_{0}.$ The bound in the statement reduces to $\gamma_{0}-\delta(1-0)=\gamma_{0},$ so the inequality is tight in this corner case.
Case 2: $\boldsymbol{m_{\mathrm{bd}}>0}$ (BCR >0>0). Now at least one boundary sample is present. By A2. its margin is $\geq-\delta$ and, because $\delta\geq\gamma_{0}$ (A3.), also $\geq\gamma_{0}-\delta$. Combining with Assumption A1. for all other points gives
@@
[…]
@@
D.3. Proof of Theorem 3.6
On-Manifold Vicinal Risk Theorem.
@@
[…]
@@
Fix $i$ temporarily. Apply A1. and law of total expectation (Grimmett and Stirzaker, 2020):
@@
$R_{\text{on}}:=\frac{1}{n}\sum_{i=1}^{n}\mathbb{E}\!\bigl[\mathbf{L}(\cdot)\mid\phi(\hat{t})\in\mathcal{M}_{c_{i}}\bigr],R_{\text{off}}:=\frac{1}{n}\sum_{i=1}^{n}\mathbb{E}\!\bigl[\mathbf{L}(\cdot)\mid\phi(\hat{t})\notin\mathcal{M}_{c_{i}}\bigr].$
Summing the displayed equality over $i$ and using A2. yields
@@
D.4. Proof of Theorem 3.7
Boundary-Coverage $\Rightarrow$ Vicinal-Risk Reduction.
@@
    Lipschitz loss. $\mathcal{J}(z_{1},y)-\mathcal{J}(z_{2},y)\leq L\|z_{1}-z_{2}\|_{2}$  (Shalev-Shwartz and Ben-David, 2014).
@@
[…]
@@
(Last three assumptions are justified in Appendix D.2.)
@@
[…]
@@
Sum (D.4) over the two synthetic subsets and divide by $m$:
[…]
@@
Choose $\eta=\delta/\gamma_{0}\;(0<\eta<1)$ and rewrite (D.4) as
@@
D.5. Proof of Theorem 3.8
Pulling Well-Aligned Nodes.
@@
[…]
@@
[…]